OpenStudy (anonymous):

Solve by substitution: {x^2+y^2=25 y-5=-x^2

OpenStudy (anonymous):

find y from second one\[y=5-x^2\]put it in the first one\[x^2+(5-x^2)^2=25\]for solving this equation let \(u=x^2\) u will get a quadratic in terms of \(u\)\[u+(5-u)^2=25\]let me know if its still unclear

OpenStudy (anonymous):

yes it is actually unclear